# Towards Accurate and Efficient 3D Object Detection for Autonomous Driving: A Mixture of Experts Computing System on Edge

### v1

@@ Abstract
This paper presents Edge-based Mixture of Experts (MoE) Collaborative Computing (EMC2), an optimal computing system designed for autonomous vehicles (AVs) that simultaneously achieves low-latency and high-accuracy 3D object detection. Unlike conventional approaches, EMC2 incorporates a scenario-aware MoE architecture specifically optimized for edge platforms. By effectively fusing LiDAR and camera data, the system leverages the complementary strengths of sparse 3D point clouds and dense 2D images to generate robust multimodal representations. To enable this, EMC2 employs an adaptive multimodal data bridge that performs multi-scale preprocessing on sensor inputs, followed by a scenario-aware routing mechanism that dynamically dispatches features to dedicated expert models based on object visibility and distance. In addition, EMC2 integrates joint hardware-software optimizations, including hardware resource utilization optimization and computational graph simplification, to ensure efficient and real-time inference on resource-constrained edge devices. Experiments on open-source benchmarks clearly show the EMC2 advancements as a end-to-end system. On the KITTI dataset, it achieves an average accuracy improvement of 3.58% and a 159.06% inference speedup compared to 15 baseline methods on Jetson platforms, with similar performance gains on the nuScenes dataset, highlighting its capability to advance reliable, real-time 3D object detection tasks for AVs.
@@ 1 Introduction
A Multimodal 3D Object Detection Solution with MoE. The proposed system achieves both high detection accuracy and inference efficiency, outperforming existing one-size-fits-all baseline models. Designed for real-world ADS applications, is optimized for deployment on resource-constrained platforms, making it more suitable for edge computing environments than generic alternatives.
Robust and Efficient Multimodal Perception through Expert Collaboration and System-Level Optimization. The proposed framework incorporates scenario-specific experts and employs an adaptive expert selection strategy to meet heterogeneous perception demands. To mitigate the training instability commonly observed in MoE-based systems, we design a hierarchical optimization scheme that combines expert-level and global backpropagation for effective collaborative learning. Furthermore, a distribution-aware resampling technique enhances robustness under real-world data imbalance. In addition to these designs, a tailored multiscale pooling module improves the efficiency of LiDAR–camera fusion, ensuring both efficient and accurate multimodal perception.
System-Level Optimization for Edge Inference. The original PyTorch-based MoE inference framework [12] suffers from inefficient management of memory and computational resources, leading to high latency on edge devices. To overcome these challenges, we propose an algorithm–hardware–software co-optimization strategy within , which incorporates efficient memory management techniques to improve L1/L2 cache hit rates and applies computational graph fusion to eliminate redundant data loading across layers. These combined optimizations notably improve inference efficiency.
SOTA Accuracy and Efficiency Combined. Experiments on the KITTI dataset demonstrate that clearly improves detection accuracy for challenging (hard-level) objects, achieving 3.28% higher accuracy for pedestrians and 6.03% for bicycles compared to baseline models. On the nuScenes dataset, further achieves a 3.9% improvement in mAP and a 1.8% boost in NDS over existing SOTA solutions. Moreover, delivers 159.06% faster inference on a Jetson platform on the KITTI benchmark, substantially enhancing real-time processing capabilities without sacrificing detection performance, and making it well-suited for edge deployment.
@@ 3.3 Scenario-Optimized Experts
Latency-Prioritized Expert (LPE). LPE utilizes 2D-CNN for object detection based on BEV projection. It first projects proposal regions in 3D voxel space, generated by the AMDB, into a BEV representation. Next, it extracts features and performs object localization and classification using a 2D CNN before mapping the detection results back into 3D space. Compared to sparse 3D convolutions, 2D convolutions offer significantly lower computational costs in outdoor ADS scenarios [15]. By leveraging this efficiency, LPE achieves high inference efficiency while maintaining reliable detection accuracy.
Versatile Efficiency Expert (VEE). VEE integrates a 3D CNN to process proposed regions and confidence scores from the AMDB. After sparse convolution, the extracted voxel features are passed through a trained decoder that generates per-voxel classifications and bounding boxes as the result. Compared to LPE, the use of 3D convolutions allows more accurate spatial structure preservation, making it more robust for scenarios with insufficient LiDAR signal.
@@ 3.6 EMC2 Computing System Optimization
Memory Optimization. To further improve memory efficiency during inference, we implement three system-level techniques. 1) Overlapping communication and computation [36]. Matrix data is partitioned into multiple segments, which are processed and transferred in parallel across processors. This overlapping of data movement and computation effectively reduces global memory access and improves cache utilization, alleviating potential memory bottlenecks. 2) Thread scheduling. We implement a staged thread management mechanism, where active threads are periodically terminated and re-initialized (Fig. 6). This approach limits unnecessary memory retention between stages, reducing system memory footprint. 3) Prefix-sum acceleration for sparse convolution. To speed up the index search for nonzero elements in sparse tensors, we apply parallel prefix-sum operations, reducing the time complexity from to by leveraging GPU parallelism, which boosts sparse convolution efficiency, and denotes the number of nonzero tensor elements.
Computational Graph Optimization. ONNX partitions the computation graph into multiple segments, dispatching each to the most suitable hardware block. Our optimization includes three stages: 1) Model pruning. Redundant parameters and operations are removed to simplify the computational graph structure. 2) Model quantization. Low-precision computations replace floating-point operations to improve hardware resource efficiency. 3) Computational graph fusion. Consecutive operators are merged to reduce memory access overhead and improve execution efficiency, which is particularly beneficial for memory-bound hardware such as CPUs, and also improves kernel launch efficiency on GPUs.
@@ 4.1 Experiment Settings and Datasets
Experiment Platform. We conduct experiments on the Jetson AGX Orin edge platform, which features limited computational resources, an NVIDIA Ampere GPU, 4MB shared L2 cache, and a hierarchical memory structure. Due to the absence of advanced sparse convolution libraries, matrix operations that exceed the L2 cache size result in notable memory exchange overhead, severely affecting computational efficiency. In addition, for comparison, we also run inference on an NVIDIA A4000 GPU.
Dataset and Sample Difficulties. We evaluate the 3D object detection performance of using two public benchmarks: KITTI [37] and nuScenes [38]. The KITTI dataset contains training samples and validation samples. According to the official KITTI difficulty levels, of the objects are labeled as “easy,” as “moderate,” and as “hard.” The nuScenes dataset consists of training scenes and test scenes, with object categories and 3D bounding box annotations for 3D detection.
Evaluation Metrics. For the KITTI dataset, detection accuracy is evaluated using the R40 3D Average Precision (AP) metric, with Intersection over Union (IoU) thresholds of for cars and for pedestrians and cyclists. The nuScenes dataset adopts the NDS metric, which combines mAP with a set of true positive error metrics. The mAP is computed based on center distance matching rather than IoU, with distance thresholds ranging from m to m depending on the object category. The true positive error metrics include mean Average Translation (mATE), Scale (mASE), Orientation (mAOE), Velocity (mAVE), and Attribute (mAAE) Errors. In addition to detection accuracy, inference latency is measured on the Nvidia Jetson AGX Orin to assess real-time deployment capability.
4.2 Implementation Details
Hierarchical Training Details. During training, we observed that allowing all experts to propagate gradients into the AMDB leads to instability. This is because the training data for different experts partially overlap, causing AMDB to be corrected multiple times for the same errors [7]. To address this, we first pre-train AMDB independently for epochs before integrating the experts. We then enable joint training with all three experts while restricting gradient propagation to AMDB solely from the APE, as APE is the most comprehensive expert that incorporates multimodality.
Data Subset Division Details. Prior to the data subset division, we conduct a statistical analysis of the confidence distribution of proposal regions generated by the aforementioned pre-trained AMDB model. The results of the two-sample Kolmogorov-Smirnov test [39] reveal a statistically notable difference in the confidence distributions of objects located within and beyond a certain distance . Specifically, the distance threshold is set to meters for the KITTI dataset and meters for the nuScenes dataset, based on empirical observations from each dataset. Accordingly, we define the target subsets for training based on the specialization of each expert as follows: 1) For the LPE, the target subset consists of scenes where all objects are within meters and labeled as “easy” or “moderate.” 2) For the VEE, the target subset includes scenes where at least one object is beyond meters but remains classified as “easy” or “moderate,” or where objects within meters are labeled as “hard.” 3) For the APE, the target subset contains scenes in which at least one object is simultaneously beyond meters and labeled as “hard,” representing the most challenging cases requiring multimodal fusion. For each expert, the auxiliary subset is uniformly sampled from the target subsets of the other experts. The number of auxiliary samples matches the number of target samples in each subset.
Adaptive Optimizer Details. We observed that when batch size is too small, the learning rate fluctuates abruptly due to the binary nature of target sample presence, leading to unstable training. Conversely, when is too large, the learning rate remains nearly constant, reducing its effect in prioritizing target samples. Through empirical evaluation, we choose a moderate batch size of to ensure the intended effect of Adaptive Optimizer while maintaining smooth learning rate adjustments.
@@ 4.3 Comparison with Existing SOTA Methods
As shown in Tab. 1 and Tab. 2, achieves SOTA performance in terms of both detection accuracy and inference latency on the KITTI and nuScenes validation sets. Furthermore, demonstrates superior performance on the Jetson AGX Orin platform, outperforming existing models across various difficulty levels.
Results on KITTI. Pedestrians and cyclists, as vulnerable road users, face higher risks in traffic [49]. However, most SOTA methods underperform in detecting these categories effectively. In Tab. 1, improves pedestrian and cyclist detection accuracy by and , respectively, on the hard setting, outperforming recent methods such as TED and LoGoNet. For car detection, achieves a robust accuracy of under the hard setting, surpassing previous multimodal approaches by up to . Importantly, demonstrates practical deployability with an average latency of ms on the Jetson AGX Orin for KITTI, representing a to speedup over existing methods that require ms to ms.
Results on NuScenes. In Tab. 2, achieves improvements of and in mAP and NDS, respectively, outperforming the latest methods FocalFormer [50] and BEVFusion [51]. Meanwhile, delivers an average inference latency of ms on the Jetson AGX Orin for nuScenes, representing a high inference speed for real-time edge deployment.
@@ 4.4 Analysis and Discussion
Benefits of Mixture-of-Experts. The balance between accuracy and efficiency demonstrated in Sec. 4.3 is achieved by dynamically selecting experts based on scenario characteristics rather than adopting a one-size-fits-all approach. This is further supported by comprehensive algorithmic and system-level optimizations tailored for edge devices with limited computational resources, such as Jetson.
Empirical Dispatcher. The dispatcher of employs a lightweight and interpretable decision-making mechanism based on fixed, empirically determined statistics, specifically distance and confidence thresholds, dynamically dispatches each scene to the most appropriate expert. Due to dataset format variations, the dispatch protocol is designed individually for each dataset. To validate this approach, we analyzed expert activations during inference on the nuScenes validation subset, as shown in Fig. 7. LPE achieves the best performance in scenes with close-range objects and high confidence scores, APE excels in distant, low-confidence scenarios, and VEE outperforms the others in all remaining cases. These observations on expert selection support the effectiveness of using empirical statistics for our dispatcher’s routing strategy.
Optimized Parameter Size. At the algorithmic level, the full contains 226M parameters in total. However, with the MoE framework, only a subset of parameters is used during each inference – averaging 87M on nuScenes and 146M on KITTI. This is substantially lower than the original and other SOTA models. At the system level, our optimization focuses not on reducing parameters, but on maximizing hardware efficiency by accelerating computation and minimizing memory access overhead.
Hardware Flexibility. The is implemented within the ONNX format, an open format supported by compilers that enables model compilation across diverse hardware platforms. As a result, is not limited to Jetson and can be deployed on other edge platforms.
4.5 Ablation Study
@@
Ablation on the Role of LPE. We further investigate the efficiency advantages of LPE compared to VEE and APE under close-range and clearly visible cases. Specifically, we assess three model variants: 1) LPE, 2) VEE, and 3) APE, all evaluated on the KITTI validation set corresponding to LPE’s target scenarios. As summarized in the first sub-table of Tab. 3, all three maintain comparable detection performance, while LPE notably reduces inference time on the Jetson platform, highlighting its effectiveness for low-latency deployment in suitable conditions.
Ablation on the Impact of VEE. We compare the performance of with and without VEE using the KITTI validation subset corresponding to VEE scenarios. The results, presented in the second sub-table of Tab. 3, demonstrate that incorporating the VEE expert leads to clear improvements in both detection accuracy and inference efficiency compared to the baseline.
Ablation on the Effect of APE. To assess the benefits of incorporating multimodal information, we conduct a comparative evaluation of VEE and APE on the full KITTI validation set. As reported in the third sub-table of Tab. 3, integrating the Image branch from the AMDB leads to a notable improvement in detection accuracy, with only a slight increase in inference time on the Jetson AGX Orin. This observation is further supported by the detection score distribution illustrated in Fig. 8.
Ablation on the Impact of Hierarchical Training (HT) and Multiscale Pooling (MP). As discussed in Sec. 3.4 and Sec. 3.5, HT directly supervises the two branches of AMDB during training and facilitates the application of MP, which enhances inference efficiency. We evaluate the combined impact of these strategies through four settings: 1) training with HT and inference with MP; 2) training with HT only; 3) inference with MP only; and 4) neither strategy used. All settings are trained for the same number of epochs and evaluated on the full KITTI validation set. The results, shown in the third sub-table of Tab. 3, indicate that although accuracy appears slightly higher when MP is not applied, combining it with HT achieves a reduction in inference time on Jetson, from ms to ms, without compromising accuracy. This is attributed to the complementary strengths of HT, which effectively learns representations at multiple levels of granularity, and MP, which leverages data compression to optimize memory utilization on resource-constrained devices. However, solely applying MP significantly reduces accuracy from to . Moreover, even employing HT alone substantially enhances accuracy from to under the same training epoch.
@@ 5 Conclusion
We introduced , a multimodal 3D object detection solution for ADS, built upon a customized MoE-based architecture. This design enables to dynamically adapt to varying traffic conditions while maintaining a practical trade-off between detection accuracy and computational efficiency. Extensive experiments on the KITTI and nuScenes datasets demonstrate that achieves state-of-the-art performance, with notable improvements under challenging conditions. Moreover, the framework delivers a inference speedup on the Jetson platform, making it highly suitable for real-time edge deployment. In the future, we will develop a fully adaptive expert selection mechanism to further improve system adaptability and edge deployment efficiency.

### v2

@@ Abstract
This paper presents Edge-based Mixture of Experts (MoE) Collaborative Computing (EMC2), an optimal computing system designed for autonomous vehicles (AVs) that simultaneously achieves low-latency and high-accuracy 3D object detection. Unlike conventional approaches, EMC2 incorporates a scenario-aware MoE architecture specifically optimized for edge platforms. By effectively fusing LiDAR and camera data, the system leverages the complementary strengths of sparse 3D point clouds and dense 2D images to generate robust multimodal representations. To enable this, EMC2 employs an adaptive multimodal data bridge that performs multi-scale preprocessing on sensor inputs, followed by a scenario-aware routing mechanism that dynamically dispatches features to dedicated expert models based on object visibility and distance. In addition, EMC2 integrates joint hardware-software optimizations, including hardware resource utilization optimization and computational graph simplification, to ensure efficient and real-time inference on resource-constrained edge devices. Experiments on open-source benchmarks clearly show the EMC2 advancements as an end-to-end system. On the KITTI dataset, it achieves an average accuracy improvement of 3.58% and a 159.06% inference speedup compared to 15 baseline methods on Jetson platforms, with similar performance gains on the nuScenes dataset, highlighting its capability to advance reliable, real-time 3D object detection tasks for AVs. The official implementation is available at https://github.com/LinshenLiu622/EMC2.
@@ 1 Introduction
1) A Multimodal 3D Object Detection Solution with MoE. The proposed system achieves both high detection accuracy and inference efficiency, outperforming existing one-size-fits-all baseline models. Designed for real-world ADS applications, is optimized for deployment on resource-constrained platforms, making it more suitable for edge computing environments than generic alternatives.
2) Robust and Efficient Multimodal Perception through Expert Collaboration and System-Level Optimization. The proposed framework incorporates scenario-specific experts and employs an adaptive expert selection strategy to meet heterogeneous perception demands. To mitigate the training instability commonly observed in MoE-based systems, we design a hierarchical optimization scheme that combines expert-level and global backpropagation for effective collaborative learning. Furthermore, a distribution-aware resampling technique enhances robustness under real-world data imbalance. In addition to these designs, a tailored multiscale pooling module improves the efficiency of LiDAR–camera fusion, ensuring both efficient and accurate multimodal perception.
3) System-Level Optimization for Edge Inference. The original PyTorch-based MoE inference framework [12] suffers from inefficient management of memory and computational resources, leading to high latency on edge devices. To overcome these challenges, we propose an algorithm–hardware–software co-optimization strategy within , which incorporates efficient memory management techniques to improve L1/L2 cache hit rates and applies computational graph fusion to eliminate redundant data loading across layers. These combined optimizations notably improve inference efficiency.
4) State-of-the-art (SOTA) Accuracy and Efficiency Combined. On the KITTI dataset, achieves 3.28% higher accuracy for pedestrians and 6.03% for bicycles under hard-level detection challenge compared to 15 baseline models. On the nuScenes dataset, it yields a 3.9% increase in mAP and a 1.8% boost in NDS. In addition, achieves a 159.06% reduction in inference latency on the Jetson platform for KITTI, supporting its suitability for real-time edge deployment.
@@ 3.3 Scenario-Optimized Experts
Latency-Prioritized Expert (LPE). LPE uses 2D CNNs for object detection based on BEV projections. It first projects proposal regions from the AMDB into BEV space, then extracts 2D features for object localization and classification. It finally maps detection results back to 3D space. Compared to sparse 3D convolutions, its use of 2D operations reduces computation in outdoor ADS scenarios [15], allowing LPE to improve inference efficiency with minimal accuracy loss.
Versatile Efficiency Expert (VEE). VEE uses a 3D CNN to process proposal regions and confidence scores from the AMDB. After sparse convolution, voxel features are decoded into per-voxel classifications and bounding boxes. Compared to LPE, 3D convolutions better preserve spatial structure, making it more robust for scenarios with insufficient LiDAR signal.
@@ 3.6 EMC2 Computing System Optimization
Memory Optimization. To further improve memory efficiency during inference, we implement three system-level techniques. 1) Overlapping communication and computation [36]. Matrix data is partitioned into segments processed and transferred in parallel, overlapping computation and communication. This reduces global memory access, improves cache usage, and mitigates memory bottlenecks. 2) Thread scheduling. We implement a staged thread management mechanism, where active threads are periodically terminated and re-initialized (Fig. 6). This approach limits unnecessary memory retention between stages, reducing system memory footprint. 3) Prefix-sum for sparse convolution. Parallel prefix-sum reduces index search time from to using GPU parallelism, where is the number of nonzero elements.
Computational Graph Optimization. ONNX partitions the computation graph into segments and assigns each to the most suitable hardware. Our optimization includes three stages: 1) Model pruning. Redundant parameters and operations are removed to simplify the graph; 2) Model quantization. Low-precision arithmetic replaces floating-point computation to improve hardware efficiency; 3) Computational graph fusion. Consecutive operators are merged to reduce memory access and improve execution, benefiting memory-bound CPUs and reducing kernel launch overhead on GPUs.
@@ 4.1 Experiment Settings and Datasets
| Method | Size | mAP | NDS | Latency (ms) |
| FocalFormer [50] | 189M | 0.6640 | 0.7090 | N/A |
| BEVFusion [51] | 156M | 0.6852 | 0.7138 | N/A |
| (ours) | 87M | 0.7241 | 0.7316 | 229.3 |
Experiment Platform. Experiments are conducted on the Jetson AGX Orin, an edge platform with limited computational resources, an NVIDIA Ampere GPU, a 4MB shared L2 cache, and hierarchical memory. For comparison, is also evaluated on an NVIDIA A4000 GPU.
Dataset and Sample Difficulties. We evaluate 3D object detection on two public benchmarks: KITTI [37] and nuScenes [38]. KITTI provides 3,712 training and 3,769 validation samples, with object difficulty levels categorized as 30% easy, 40% moderate, and 30% hard. nuScenes contains 850 training and 150 test scenes, covering 10 object categories with 3D bounding box annotations. Before splitting the data, we analyze the confidence distribution of proposal regions generated by the pre-trained AMDB. A two-sample Kolmogorov–Smirnov test reveals a statistically significant difference between objects within and beyond a distance threshold , which is empirically set to meters for KITTI and meters for nuScenes. The training subsets for each expert on the KITTI and nuScenes datasets are defined as follows: includes scenes where all objects are within meters and labeled as “easy” or “moderate”; includes scenes with at least one object beyond meters labeled as “easy” or “moderate,” or any object within meters labeled as “hard”; and includes scenes where at least one object is both beyond meters and labeled as “hard,” requiring multimodal processing. For each expert, the auxiliary subset is uniformly sampled from the other experts’ target subsets, with its size equal to that of the corresponding target subset.
Evaluation Metrics. KITTI uses R40 3D Average Precision (AP) with IoU thresholds of 0.7 for cars and 0.5 for pedestrians and cyclists. nuScenes adopts the NDS metric, combining mAP—computed via center distance matching (0.5 m to 4.0 m thresholds)—and five error terms: mean Average Translation (mATE), Scale (mASE), Orientation (mAOE), Velocity (mAVE), and Attribute (mAAE) Errors. Inference latency on Jetson AGX Orin is also reported to assess real-time performance.
Hierarchical Training Details. Allowing all experts to propagate gradients to AMDB causes instability due to overlapping training data, leading to repeated updates. To address this, AMDB is first pre-trained for 20 epochs, followed by joint training where gradients are propagated only from APE, which receives full multimodal input.
@@ 4.2 Comparison with Existing SOTA Methods
As shown in Tab. 1 and Tab. 2, achieves SOTA accuracy and latency on KITTI and nuScenes. On the Jetson AGX Orin platform, it also outperforms existing baseline models across all difficulty levels.
Results on KITTI. Pedestrians and cyclists face higher traffic risks [49], which remain challenging for most SOTA methods. improves hard-set detection by and for pedestrians and cyclists, respectively, outperforming TED and LoGoNet (Tab. 1). For cars, it achieves on the hard set, exceeding prior multimodal methods by up to . It runs at ms on Jetson, – faster than methods requiring – ms.
Results on nuScenes. improves mAP and NDS by and , respectively, outperforming recent methods FocalFormer [50] and BEVFusion [51] (Tab. 2). It runs at ms on Jetson AGX Orin, enabling real-time use on edge devices.
@@ 4.3 Analysis and Discussion
Benefits of Mixture-of-Experts. The balance between accuracy and efficiency observed in Sec. 4.2 is achieved through dynamic expert selection based on scenario characteristics, combined with algorithmic and system-level optimizations designed for edge devices, such as Jetson.
Empirical Dispatcher. The dispatcher uses fixed distance and confidence thresholds to dispatch scenes to experts. The protocol is adapted per dataset format. This strategy is validated via expert activations on the nuScenes validation set (Fig. 7): LPE performs best in close-range, high-confidence scenes; APE in distant, low-confidence cases; and VEE in all others, supporting the effectiveness of empirical routing.
Optimized Parameter Size. has 226M parameters, but its MoE framework activates only a subset during inference, averaging 87M on nuScenes and 146M on KITTI, which is fewer than those used by the full model and other SOTA baselines. System-level optimizations instead prioritize computational throughput and memory efficiency over parameter reduction.
Hardware Flexibility. The is implemented in the ONNX format, an open format supported by various compilers and enables model compilation across diverse hardware platforms. As a result, is not limited to Jetson and can be deployed on other edge platforms.
@@
4.4 Ablation Study
Ablation on the Roles of Three Experts. We evaluate each expert on its target scenarios using the KITTI validation set. As shown in Tab. 3, LPE, VEE, and APE achieve comparable accuracy in LPE scenes, but LPE has the lowest inference latency on Jetson. In VEE scenarios, incorporating VEE improves both accuracy and latency over the baseline. For APE, which incorporates image features, accuracy improves further with only a minor latency increase, supported by the score distribution in Fig. 8.
Ablation on Hierarchical Training (HT) and Multiscale Pooling (MP). We evaluate four combinations of HT and MP on the full KITTI validation set: both applied, HT only, MP only, and neither. Tab. 3 shows that combining HT and MP reduces inference time on Jetson from ms to ms without accuracy loss. For the car class, MP alone reduces accuracy from to , while HT alone improves it from to . HT facilitates multigranular feature learning, while MP enhances memory efficiency via compression.
@@ 5 Conclusion
We present , a multimodal 3D object detection solution for ADS based on a customized MoE architecture. This design enables adaptive expert selection across diverse traffic scenarios, balancing detection accuracy and efficiency. Experiments on KITTI and nuScenes show strong performance, particularly under challenging conditions. On the Jetson platform, achieves a speedup in inference, supporting real-time edge deployment. Future work will explore fully adaptive expert selection to enhance system responsiveness and resource efficiency.